OpenStudy (anonymous):

Media Services charges $20 for a phone and $30 for its economy plan. The equation c = 30t + 20 describes the total cost, c, of operating a Media Services phone for t months. That is the total cost for t months of service? If customer has only $90 available, how many months of service can they receive?

OpenStudy (anonymous):

So in this case you are given "C", and need to solve for "t". 90=30t+20 70=30t 70/30=t t=2.33(repeating) So they can pay for 2 full months, as the ".333" isn't a full month.

OpenStudy (anonymous):

Yea.

OpenStudy (anonymous):

I am lost.

OpenStudy (anonymous):

Would it be 2 1/3?

OpenStudy (anonymous):

That is the answer to the EQUATION yeah, but to the actual QUESTION, they could only pay for 2 months, because you only consider the highest whole number, which here is 2.

OpenStudy (anonymous):

2 1/2

OpenStudy (anonymous):

2/1/2??????

OpenStudy (anonymous):

The answer to the question is 2. With the costs and how much they can pay, they are able to pay for 2 full months.

OpenStudy (anonymous):

oh crud. Thank you. The total cost for 3 months of service is???

OpenStudy (anonymous):

$3?

OpenStudy (anonymous):

if you wanted 3 months of service you would set the equation up like this: c=30(3)+20 c=90+30 c=120 it would be 120$ for 3 months of service

OpenStudy (anonymous):

hmm then $3 would not be right.

OpenStudy (anonymous):

$90

OpenStudy (anonymous):

$9

OpenStudy (anonymous):

Oh crap

OpenStudy (anonymous):

$120

OpenStudy (anonymous):

Yeah, it would be 120$ for 3 months

OpenStudy (anonymous):

gosh, I am a moron. Okay so if a customer has just $90 how many month of service can they receive?

OpenStudy (anonymous):

2 months

OpenStudy (anonymous):

2 1/2 months

OpenStudy (anonymous):

2 months?

OpenStudy (anonymous):

Yeah, because you can't pay for half a month of service, you have to round down to 2.